# Distance-Domain Degrees of Freedom in Near-Field Region

### v3

@@ VII Conclusion
In this paper, we investigated the distance-domain DoF by considering a MIMO channel where the source (represents base station) equips with an Tx array with arbitrary shape and the observation region (represents users) is Rx linear array whose elements are aligned at the same angle but lie at different distances from the Tx array. In an ideal scenario where the Tx array is single-piece and the Rx array is on the broadside of the Tx array, we consider both the arrays having continuous aperture with infinite elements and infinitesimal spacing and model the channel as an integral operator with a non-convolution kernel. We further transform it into the integral operator with a Hermitian convolution kernel, which enables eigenvalue analysis via the Fourier transform. This formulation allowed us to derive a closed-form expression for the distance-domain DoF, demonstrating that unlike the angular-domain DoF, the distance-domain DoF does not exhibit linear relationship with aperture’s area and mainly determined by the extreme boundaries of the Tx and Rx arrays rather than its detailed shape.
@@
By performing the variable substitution $t=1/r$ with the factor $dr=dt/t^{2}$) in (63), we obtain
$\lambda_{k}\,e_{k}\!\left(\frac{1}{t}\right)$ $=\int_{r^{-1}_{\max}}^{r^{-1}_{\min}}{v}\!\left(\frac{1}{t},\frac{1}{t^{\prime}}\right)\,e_{k}\!\left(\frac{1}{t^{\prime}}\right)\left(\frac{1}{t^{\prime 2}}\,dt^{\prime}\right).$
Multiplying both sides by $1/t$ and let $f_{k}(t)=e_{k}(1/t)/t$:
$\lambda_{k}\,f_{k}(t)$ $=\frac{1}{t}\,\lambda_{k}\,e_{k}\!\left(\frac{1}{t}\right)$
$=\frac{1}{t}\int_{r^{-1}_{\max}}^{r^{-1}_{\min}}{v}\!\left(\frac{1}{t},\frac{1}{t^{\prime}}\right)\,e_{k}\!\left(\frac{1}{t^{\prime}}\right)\left(\frac{1}{t^{\prime 2}}\,dt^{\prime}\right)$
$=\int_{r^{-1}_{\max}}^{r^{-1}_{\min}}\left[\frac{1}{t\,t^{\prime}}\,{v}\!\left(\frac{1}{t},\frac{1}{t^{\prime}}\right)\right]\left[\frac{1}{t^{\prime}}e_{k}\!\left(\frac{1}{t^{\prime}}\right)\right]\,dt^{\prime}$
$=\int_{r^{-1}_{\max}}^{r^{-1}_{\min}}g(t,t^{\prime})\,f_{k}(t^{\prime})\,dt^{\prime},$ (64)
@@
$g(t,t^{\prime})=\frac{1}{t\,t^{\prime}}\,{v}\!\left(\frac{1}{t},\frac{1}{t^{\prime}}\right).$ (65)
Thus, $f_{k}(t)$ and $\lambda_{k}$ satisfy the eigenvalue equation for the operator $\mathcal{G}$ with kernel $g(t,t^{\prime})$, which completes the proof.
@@
$\tilde{g}(\xi)=\int_{-\infty}^{\infty}g(t)e^{-j2\pi\xi\frac{t}{T}}dt,$
where $T=r^{-1}_{\min}-r^{-1}_{\max}$. We express $g(t)=g_{0}(t)w(t)$ where $w(t)=\mathbf{1}_{\{|t|\leq T\}}$ is the window function with duration $2T$ and $g_{0}(t)$ is given by:
$g_{0}(t)=\iint_{\mathbf{p}\in\mathscr{P}}\exp\!\Bigl(-j\frac{\pi}{\lambda}\|\mathbf{p}\|^{2}\,t\Bigr)d\mathbf{p}.$
while the Fourier transform of $w(t)$ is simply $\tilde{w}(\xi)=2\,\mathrm{sinc}(2\xi)$ with $\mathrm{sinc}(x)=\frac{\sin(\pi x)}{\pi x}$. In addition, interchanging the order of integration shows that the scaled Fourier transform of $g_{0}(t)$ is
$\tilde{g}_{0}(\xi)$ $=\int_{-\infty}^{\infty}\left[\iint_{\mathbf{p}\in\mathscr{P}}e^{-j\frac{\pi}{\lambda}\|\mathbf{p}\|^{2}\,t}d\mathbf{p}\right]e^{-j2\pi\xi\frac{t}{T}}dt$
$=\iint_{\mathbf{p}\in\mathscr{P}}\left[\int_{-\infty}^{\infty}e^{-j2\pi\left(\frac{\|\mathbf{p}\|^{2}}{2\lambda}+\frac{\xi}{T}\right)\,t}dt\right]d\mathbf{p}$
$=T\iint_{\mathbf{p}\in\mathscr{P}}\delta\!\Bigl(\xi+\frac{\|\mathbf{p}\|^{2}T}{2\lambda}\Bigr)d\mathbf{p}$ (66)
By the convolution theorem, the Fourier transform of $g(t)$ is the convolution of the Fourier transforms of the two factors. And by ignoring the constant scalar coefficients, we obtain:

### v4

@@ VIII Conclusion
In this paper, we investigated the distance-domain DoF by considering a MIMO channel where the source (represents base station) equips with an Tx array with arbitrary shape and the observation region (represents users) is Rx linear array whose elements are aligned at the same angle but lie at different distances from the Tx array. In an ideal scenario where the Tx array is single-piece and the Rx array is on the broadside of the Tx array, we consider both the arrays having continuous aperture with infinite elements and infinitesimal spacing and model the channel as an integral operator with a non-convolution kernel. We further transform it into the integral operator with a Hermitian convolution kernel, which enables eigenvalue analysis via the Fourier transform. This formulation allowed us to derive a closed-form expression for the distance-domain DoF. Our results demonstrate that, unlike the angular-domain DoF, the distance-domain DoF does not scale linearly with the aperture’s area. Instead, it is primarily determined by the extreme radial boundaries of the Tx and Rx arrays, regardless of their detailed interior shapes.
@@
By performing the variable substitution $\zeta=1/r$ with the factor $dr=d\zeta/\zeta^{2}$) in (55), we obtain
$\lambda_{k}\,e_{k}\!\left(\frac{1}{\zeta}\right)$ $=\int_{r^{-1}_{\max}}^{r^{-1}_{\min}}{v}\!\left(\frac{1}{\zeta},\frac{1}{\zeta^{\prime}}\right)\,e_{k}\!\left(\frac{1}{\zeta^{\prime}}\right)\left(\frac{1}{\zeta^{\prime 2}}\,d\zeta^{\prime}\right).$
Multiplying both sides by $1/\zeta$ and let $f_{k}(\zeta)=e_{k}(1/\zeta)/\zeta$:
$\lambda_{k}\,f_{k}(\zeta)$ $=\frac{1}{\zeta}\,\lambda_{k}\,e_{k}\!\left(\frac{1}{\zeta}\right)$
$=\frac{1}{\zeta}\int_{r^{-1}_{\max}}^{r^{-1}_{\min}}{v}\!\left(\frac{1}{\zeta},\frac{1}{\zeta^{\prime}}\right)\,e_{k}\!\left(\frac{1}{\zeta^{\prime}}\right)\left(\frac{1}{\zeta^{\prime 2}}\,d\zeta^{\prime}\right)$
$=\int_{r^{-1}_{\max}}^{r^{-1}_{\min}}\left[\frac{1}{\zeta\,\zeta^{\prime}}\,{v}\!\left(\frac{1}{\zeta},\frac{1}{\zeta^{\prime}}\right)\right]\left[\frac{1}{\zeta^{\prime}}e_{k}\!\left(\frac{1}{\zeta^{\prime}}\right)\right]\,d\zeta^{\prime}$
$=\int_{r^{-1}_{\max}}^{r^{-1}_{\min}}g(\zeta,\zeta^{\prime})\,f_{k}(\zeta^{\prime})\,d\zeta^{\prime},$ (56)
@@
$g(\zeta,\zeta^{\prime})=\frac{1}{\zeta\,\zeta^{\prime}}\,{v}\!\left(\frac{1}{\zeta},\frac{1}{\zeta^{\prime}}\right).$ (57)
Thus, $f_{k}(\zeta)$ and $\lambda_{k}$ satisfy the eigenvalue equation for the operator $\mathcal{G}$ with kernel $g(\zeta,\zeta^{\prime})$, which completes the proof.
@@
$\tilde{g}(\xi)=\int_{-\infty}^{\infty}g(\zeta)e^{-j2\pi\xi\frac{\zeta}{L_{\zeta}}}d\zeta,$
where $L_{\zeta}=r^{-1}_{\min}-r^{-1}_{\max}$. We express $g(\zeta)=g_{0}(\zeta)w(\zeta)$ where $w(\zeta)=\mathbf{1}_{\{|\zeta|\leq L_{\zeta}\}}$ is the window function with duration $2T$ and $g_{0}(\zeta)$ is given by:
$g_{0}(\zeta)=\iint_{\mathbf{p}\in\mathscr{P}}\exp\!\Bigl(-j\frac{\pi}{\lambda}\|\mathbf{p}\|^{2}\,\zeta\Bigr)d\mathbf{p}.$
while the Fourier transform of $w(\zeta)$ is simply $\tilde{w}(\xi)=2\,\mathrm{sinc}(2\xi)$ with $\mathrm{sinc}(x)=\frac{\sin(\pi x)}{\pi x}$. In addition, interchanging the order of integration shows that the scaled Fourier transform of $g_{0}(\zeta)$ is
$\tilde{g}_{0}(\xi)$ $=\int_{-\infty}^{\infty}\left[\iint_{\mathbf{p}\in\mathscr{P}}e^{-j\frac{\pi}{\lambda}\|\mathbf{p}\|^{2}\,\zeta}d\mathbf{p}\right]e^{-j2\pi\xi\frac{\zeta}{L_{\zeta}}}d\zeta$
$=\iint_{\mathbf{p}\in\mathscr{P}}\left[\int_{-\infty}^{\infty}e^{-j2\pi\left(\frac{\|\mathbf{p}\|^{2}}{2\lambda}+\frac{\xi}{L_{\zeta}}\right)\,\zeta}d\zeta\right]d\mathbf{p}$
$=L_{\zeta}\iint_{\mathbf{p}\in\mathscr{P}}\delta\!\Bigl(\xi+\frac{\|\mathbf{p}\|^{2}T}{2\lambda}\Bigr)d\mathbf{p}$ (58)
By the convolution theorem, the Fourier transform of $g(\zeta)$ is the convolution of the Fourier transforms of the two factors. And by ignoring the constant scalar coefficients, we obtain: